March 22, 2017, Production & Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Wednesday March 22, 2017 at 2:00 pm EDT. For those people at Jefferson Lab, the meeting will be in room F326-327.

Meeting Connections

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Reconstruction & Analysis Studies Status


  1. Announcements
  2. Data Production
  3. Analysis Focus
  4. This Week's Studies
  5. Any other studies
  6. Upcoming Study Schedule

This Week's Topics

  • Talks: Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • Talks should be limited to 10 + 5 minutes.

Data Production

  1. Monitoring Update --- Thomas Britton
  2. Calibration Update --- Sean Dobbs
  3. Preparing for the first Spring 2017 production launch
  4. Processing Update --- Alex Austregesilo

Analysis Focus: Total Cross Sections

  1. J/Ψ, φ Lubomir
  2. η, ρ, ω, φ ω cross section update -- Simon
  3. φ --- Thomas
  4. Any others?

Studies: Other

Updates on experiment, simulation, and the comparison between the two.

  1. Track / Hit Matching: BCAL, FCAL, TOF, SC --- Paul Mattione
  2. Efficiencies: TOF, SC
    1. TOF Stability during Spring 2017: ADC Peak Amplitude during the run period
  3. Means & Resolutions (time, energy, dE/dx): Tracking, BCAL, FCAL, SC, TOF
  4. Uncertainties: PID (BCAL, FCAL, TOF, dE/dx), Kinfit (BCAL, FCAL, tracking)
  5. Channel/Analysis Studies: Branching ratios, cross sections, SDMEs, beam asymmetries
  6. Other reconstruction/analysis issues

Data Production Milestones

  1. Fri. April 7 - Finalize calibrations/software for low-rate running. Run final tests.
  2. Fri. April 14 - Start production for low-rate running.
  3. Fri. April 21 - Finalize calibrations/software for high-rate running. Run final tests.
  4. Fri. April 28 - Start production for high-rate running.

Upcoming Study Schedule

  • Updates on experiment, simulation, and the comparison between the two.
  • Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • All talks should be limited to 10 + 5 minutes.
  • Next Week: Beamline & Triggering
  1. Flux --- Justin
  2. Beam energy --- Beamline Group
  3. Polarization (TPOL & lineshape) --- Beamline Group
  4. Trigger emulation --- Alex Somov
  5. Triggering efficiency --- Alex Somov
  • Week after next: Hit Efficiencies (i.e. Is the detector working?)
  1. CDC Hit Efficiencies --- Naomi Jarvis
  2. FDC Hit Efficiencies --- Alex Austregesilo
  3. BCAL Hit Efficiencies --- Elton Smith
  4. FCAL Hit Efficiencies --- Jon Zarling
  • The following week: Alignment & Track/Shower Efficiencies

(i.e. Is the reconstruction working?) Updates on experiment, simulation, and the comparison between the two.

  1. Drift chamber alignment --- Mike Staib
  2. Tracking Efficiencies --- Paul Mattione
  3. BCAL Shower Efficiencies --- Elton Smith
  4. FCAL Shower Efficiencies --- Jon Zarling


Data Production

  1. Sean reviewed the current status of calibrations. Initial "acceptable" calibrations have been done for the low-rate running, except for a few trigger issues tracked on this wiki page. He is finalizing the calibrations for the high rate running. Higher level calibrations are being steadily completed; he reminded that these should be checked to see if they are stable over the run. Also, the skim jobs are >90% finished.
    • Please let Sean know if any other reconstruction issues need to be fixed before the reconstruction launch, in order to generate physics-quality data.
    • Sean is also organizing help for detector experts to check the results of the monitoring launches. All levels of experience are welcome! Send him an email if you are interested. He also encouraged people to look at results of exclusive reactions on the REST files generated by the monitoring launches.
  2. Justin prepared wiki pages describing the Spring 2017 Dataset and the corresponding Trigger Conditions, and reviewed their contents.
    • It was pointed out that the FCAL HV changes should lead to a difference in the reconstructed energy resolutions, and the TAGM changes shouldn't lead to any changes if the standard software cuts in CCDB are used.
  3. Alex A. launched monitoring jobs over the low-rate data this morning after getting the go-ahead from Sean. The plan is to launch the high-rate runs tomorrow.
    • Sean laid out proposed milestones for finalizing calibrations and producing data in an email on Friday; they are listed in the agenda. Mike S. pointed out that once the tracking alignment is finalized, we will want to do a final production of the Spring 2016 data. Depending on when this is finalized, we will work that production into the schedule.

Analysis Focus: Total Cross Sections

  1. Simon presented a look at omega cross sections using 5 files from one amorphous run, and using the energy of the reconstructed omega -> pi+pi-pi0 decay instead of any tagger information. The resulting cross sections were generally within ~10% of the expected cross section in the PS energy range. A next step will be to look at the REST files from the new monitoring launch where the tagger is properly calibrated.
  2. Thomas is making progress on the simulation front; this was discussed in some detail at this morning's Offline Software meeting.

Studies: Other

  1. Paul showed the updates on the latest iteration of the track/hit matching cuts and how they look in the standard monitoring histograms. There have been many improvements and some features that are not entirely understood, however more detailed studies should be done by the detector groups to optimize these cuts and determine the correct efficiencies.
  2. Beni showed some results on the stability of the TOF ADC peak amplitudes over the last run. The distributions were smoothly varying, with well-understood features, compared to the Spring 2016 run period, when they were varying wildly from run to run.